OpenStudy (anonymous):

if x(cannont equal)5, what is the answer to the Expression below in Simplest form? 100-4x^/5-x

OpenStudy (anonymous):

factor the numerator and then cancel

OpenStudy (anonymous):

what do you mean by factor

OpenStudy (anonymous):

is it 4x^2?

OpenStudy (asnaseer):

assuming the numerator is \(100-4x^2\), then notice that 100 and 4 are both multiples of 4, so you can factor a 4 out to get:\[100-4x^2=4(25-x^2)\]then use the following rule:\[a^2-b^2=(a+b)(a-b)\]to simplify the numerator further. you should then be able to cancel out a common term in the numerator and denominator as satellite73 suggested.

OpenStudy (anonymous):

FOILS? (a+b)(a-b)?

OpenStudy (asnaseer):

yes - this is known as the difference of two squares

OpenStudy (asnaseer):

look carefully at \(25-x^2\) and see if you can re-write as the difference of two squares.

OpenStudy (asnaseer):

so 25 is something squared - what is that something?

OpenStudy (anonymous):

5

OpenStudy (asnaseer):

correct, therefore:\[25-x^2=5^2-x^2\]now use the factorisation for the difference of two squares.

OpenStudy (asnaseer):

i.e. use:\[a^2-b^2=(a+b)(a-b)\]where for this question a=5 b=x

OpenStudy (asnaseer):

do you understand ZSpence?

OpenStudy (anonymous):

yes ty

OpenStudy (asnaseer):

ok, good. so you can solve from here I assume?

OpenStudy (anonymous):

yes ty

OpenStudy (asnaseer):

yw